# Non-unital noise in a superconducting quantum computer as a computational resource for reservoir computing

### v2

@@ Acknowledgements
Supplementary material to: ”Non-unital noise as a computational resource for quantum reservoir computing”
I Quantum reservoir computing and fading memory
@@
Definition I.1 (Fading memory).
@@
II A geometrical interpretation of non-unital channels
@@
In particular, combining Theorem II.2 and Theorem II.3, we can deduce that a channel with exactly two pure outputs is unital if and only if they are antipodal on the Bloch sphere (thus, are orthogonal in the Hilbert space). Fig. 5 in the Main text shows the action of amplitude damping, depolarizing, phase damping, and fake-OSAKA channel. As expected, amplitude damping is the sole non-unital ideal noise channel. Moreover, we can observe that the fake-OSAKA channel is non-unital.
III Contractivity of noise channels
@@
IV Universal approximation property of our gate-based echo state network
@@
Lemma IV.2.
The space of bounded time-dependent sequences $\mathcal{I}$ equipped with the distance $d_{\omega}$ defined in Def. I.1 is a compact metric space.

### v3

@@ Acknowledgements
Appendix A Supplementary Material
Appendix B Quantum reservoir computing and fading memory
@@
Definition 1 (Fading memory).
@@
Appendix C A geometrical interpretation of non-unital channels
@@
In particular, combining Theorem 3 and Theorem 4, we can deduce that a channel with exactly two pure outputs is unital if and only if they are antipodal on the Bloch sphere (thus, are orthogonal in the Hilbert space). Fig. 5 in the Main text shows the action of amplitude damping, depolarizing, phase damping, and fake-OSAKA channel. As expected, amplitude damping is the sole non-unital ideal noise channel. Moreover, we can observe that the fake-OSAKA channel is non-unital.
Appendix D Contractivity of noise channels
@@
Appendix E Universal approximation property of our gate-based echo state network
@@
Lemma 1.
The space of bounded time-dependent sequences $\mathcal{I}$ equipped with the distance $d_{\omega}$ defined in Def. 1 is a compact metric space.